OpenStudy (anonymous):

What are all the roots of x^3+x^2+4x+4=0 ?

OpenStudy (anonymous):

x^3+x^2+4x+4=0 x^3+4x+x^2+4=0 x(x^2+4)+1(x^2+4)=0 (x+1)(x^2+4)=0

OpenStudy (anonymous):

Now, can u?

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

it makes sense now